OpenStudy (anonymous):

Solve on the interval [o,2pi) 1+ Sin(Theta)=1/2

OpenStudy (aravindg):

simple take 1 to other side

OpenStudy (aravindg):

sin theta=-1/2

OpenStudy (aravindg):

then u can easily solve for angles where sin theta =-1/2

OpenStudy (anonymous):

but the Answers are A. 7pi/6,11pi/6 B.pi/3,5pi/3 C. 2pi/3,4pi/3 D.pi/6,5pi/6

OpenStudy (aravindg):

cant u find dout that :P

OpenStudy (aravindg):

one angle will be\[\huge \pi+\frac{\pi}{6}=\frac{7\pi}{6}\]

OpenStudy (aravindg):

cant u find the other?

OpenStudy (anonymous):

sin is negative in the 3rd and 4th quadrants

OpenStudy (anonymous):

i really don't get it tho.

OpenStudy (aravindg):

other is 330 degrees

OpenStudy (aravindg):

convert into radians :P

OpenStudy (anonymous):

:/

OpenStudy (aravindg):

gt it?

OpenStudy (anonymous):

C or B?

OpenStudy (aravindg):

ITS A!!!

OpenStudy (aravindg):

THE OTHER ANGLE IS 11 PI/6

OpenStudy (aravindg):

no rewards ?:P

OpenStudy (aravindg):

@yessica92

OpenStudy (anonymous):

OK Thanks!!

OpenStudy (aravindg):

u didnt giv me best answer